OpenStudy (anonymous):

Find the distance between the points (-3,7) and (4,12).

OpenStudy (anonymous):

distance from-3 to 4 is 7 distance from 7 to 12 is 5 pythagoras gives \[\sqrt{7^2+5^2}\] \[\sqrt{49+25}\] \[\sqrt{74}\]

OpenStudy (anonymous):

x2-x1 + y2-y1